OpenStudy (anonymous):

Multiply and simplify by factoring. Assume that all expressions under radicals represent nonnegative numbers. ^3√y^13 ^3√81y^14 I came up with ^3√9y^1

OpenStudy (saifoo.khan):

\[\Large \sqrt[3]{y^{13}} * \sqrt[3]{81 y^{14}} \to \sqrt[3]{81 y^{27}}\]

OpenStudy (saifoo.khan):

\[\Large 3 y^9 \sqrt [3]{3}\]

OpenStudy (anonymous):

Thanks!